OpenStudy (anonymous):

has a slope of -1/2 and has a x intercept of -2 ? write in slope intercept form.

OpenStudy (anonymous):

@sorour help?

OpenStudy (anonymous):

the point is (-2,0) \[y - y _{1} =m \times (x - x _{1})\]

OpenStudy (anonymous):

m=-1/2 \[y _{1} =0\] \[x _{1} =-2\]

OpenStudy (anonymous):

@broderick365 can u get it?

OpenStudy (anonymous):

0= -1/2 x (x-2)?? @sorour

OpenStudy (anonymous):

y=-1/2*(x-2)

OpenStudy (anonymous):

sorry y=-1/2*(x-(-2))

OpenStudy (anonymous):

is that the answer @sorour

OpenStudy (anonymous):

yes,the fional answer is: y=-1/2x-1

OpenStudy (anonymous):

tanx for a medal :)